OpenStudy (anonymous):

Given , complete the square and convert the equation to y = a(x − h)2 + k.

OpenStudy (anonymous):

y=-3x^2+30x+7

OpenStudy (anonymous):

y=-3(x^-10x+25)+75+7 y=-3(x-5)^2+82

OpenStudy (anonymous):

THANKS BRO!

OpenStudy (anonymous):

:)